use inductive reasoning to predict the most probable next number in the list 7,9,5,7,3,5,1,3,-1,?
Answer by greenestamps(13200)



7+2=9; 9-4=5
5+2=7; 7-4=3
3+2=5; 5-4=1
1+2=3; 3-4=-1

The apparent pattern is "add 2, subtract 4, repeat". So the most probable next number in the sequence is

-1+2=1

Alternatively, separate the given sequence into two spliced sequences:
7   5   3   1  -1
  9   7   5   3   x

In each sub-sequence, the pattern is to repeatedly subtract 2. That means the next number (in the second sub-sequence) is 1.